OpenStudy (jessie13):

Will Medal!Answer ASAP!! 1: All of the following expressions are equivalent except _____. 7 - x x - 7 - x + 7 7 + (- x) 2: Solve -16x = -80. -96 -64 -5 5 3: The solution to -45 = n/5 is _____. -9 9 225 -225

OpenStudy (jessie13):

@phi

OpenStudy (anonymous):

1.c

OpenStudy (anonymous):

2.d

OpenStudy (jessie13):

do u know the last one? and what grade are u in?

OpenStudy (anonymous):

I'm in 9th but no I don't know

OpenStudy (jessie13):

oh okay but thanks any way?

OpenStudy (phi):

Can you answer #1 ? do any look like they are the same expression (written in different ways) ?

OpenStudy (jessie13):

D

OpenStudy (phi):

you should know that you can change the order of when adding x+ -7 is the same as -7+ x and putting in parens suchas (-7) + x does not change it.

OpenStudy (phi):

all the choices have +7 and -x being added except for one of the choices that adds -7 and +x

OpenStudy (jessie13):

ok

OpenStudy (jessie13):

but did @cool4u2001 get the answers right?

OpenStudy (phi):

in Q1, which choice has a -7 in it ?

OpenStudy (jessie13):

b

OpenStudy (phi):

notice none of the other choices have -7 b is the "different" one

OpenStudy (jessie13):

oh! so b is right?

OpenStudy (phi):

yes, b for Q1 for Q2, what is -16* 5 ?

OpenStudy (jessie13):

-80

OpenStudy (phi):

so -16*5 = -80 -16*x = -80 x must be what ?

OpenStudy (jessie13):

5

OpenStudy (phi):

so d is the answer for Q2 btw, you find x by starting with -16 x = -80 divide both sides by -16 x= -80/-16 = 5

OpenStudy (phi):

Q3 does not make sense. It looks like something was left out.

OpenStudy (jessie13):

i fixed it!

OpenStudy (jessie13):

all i know about it is that its not A

OpenStudy (jessie13):

@phi

OpenStudy (jessie13):

@phi

OpenStudy (phi):

multiply both sides by 5

OpenStudy (jessie13):

what do u mean?

OpenStudy (phi):

\[ -45\cdot 5 =\frac{ n}{5} \cdot 5 \]

OpenStudy (phi):

on the right side 5/5 simplifies to 1 and 1*n is just n you get -45*5 = n

OpenStudy (jessie13):

-225?

OpenStudy (phi):

yes

OpenStudy (jessie13):

okay thanks for helping me!!!!

OpenStudy (phi):

yw

OpenStudy (jessie13):

thank so much i got 100%!!!!

OpenStudy (jessie13):

@phi